PLEASE HELP
I need this within 10 minutws for the math exam.
Find the shortest distance from the point P(1,0) to the line x-2y+4=0 to the nearest tenth.
A. 5.5
B. 9.2
C. 6.3
D. 2.2

Answer by venugopalramana(3286)

I need this within 10 minutws for the math exam.
Find the shortest distance from the point P(1,0)=(X1,Y1)SAY to the line x-2y+4=0...AX+BY+C=0 SAY to the nearest tenth.
FORMULA IS [AX1+BY1+C]/|SQRT(A^2+B^2)|
=(1-2*0+4)/SQRT(1^2+2^2)=5/SQRT(5)=SQRT(5)
D=2.2 IS THE ANSWER

A. 5.5
B. 9.2
C. 6.3
D. 2.2